carmelle:

What are the x-intercepts of the graph of f(x)? f(x) = 2x^2 − x − 10 *I have to solve for it and not use desmos* ._.

snowflake0531:

Plug and chug (2x +/- _ )(x +/- _) You know that we have factors of 1, 10 or 2, 5 Obviously it can't be 1, 10, or else no matter what order we put it in, we won't get -x So 2 and 5, which one's positive, which one's negative, which one gets multiplied with x, which one gets multiplied with 2x?

carmelle:

:( im confused-

snowflake0531:

By the way, a hint 2x times 2 is 4x, while 5x alone is 5x, adn 4x - 5x is -x So whichone is negative, which one is positive which one is multiplied by 2

snowflake0531:

...

snowflake0531:

(2x +/- __)(x +/- __) you do get how that is made, right?

carmelle:

uh- is it factoring

snowflake0531:

2x times x is 2x^2, that's why we have 2x and x

carmelle:

ah ok, I understand

snowflake0531:

Okay so now try them all in, you know for sure we use 2 and 5, right?

carmelle:

yea? bc 2 x 5 = 10

snowflake0531:

Yes so then it's (2x +/- 2or5)(x +/- 2or5)

snowflake0531:

Since I have no idea how to coherently explain this to you again, i suggest trying each of them -2, 5 2, -5 -5, 2 -2, 5 o-O

Angle:

want me to try explaining?

snowflake0531:

yes angle explain o-o

Angle:

if you have been taught the FOIL method before, this is just that idea but backward > yes it is factoring

Angle:

you are looking for two numbers that: multiply to get -20 [because the first 2 and the last -10 multiplied is -20] and add up to -1 [because the middle coefficient is -1 times the middle x]

Angle:

I will rephrase: two numbers that multiply to get -20 but also add up to -1

Angle:

usually people use "guess and check" to get their numbers 1, 20 2, 10 4, 5 all multiply to get 20 but only 4 - 5 = -1 this means we want to use 4 and -5 as our numbers

carmelle:

ah ok, I understand

Angle:

awesome now, what I am going to do next is called factoring by grouping It uses the idea that 4x - 5x = -x which means we can go the other way and split up the -x in the middle -x = 4x - 5x means 2x^2 - x - 10 turns into 2x^2 + 4x - 5x - 10

Angle:

If I have 2x^2 + 4x - 5x - 10 this can also look like (2x^2 + 4x) + (-5x + -10) can you try factoring out the greatest common factor out of each pair?

carmelle:

uh.. 2x(x+2) and -5(x+2)

Angle:

amazing!

Angle:

then notice how they have (x+2) in common you can factor that to the front as well

carmelle:

:0 I thought I was wrong- ty

carmelle:

so would it be... (x+2)(2x-5)

Angle:

awesome! (x+2)(2x-5) = 0 do you know how to solve for x in this?

carmelle:

umm, I dont think so 0-0

Angle:

you take each of the stuffs in the parenthesis and set it equal to 0 x+2 = 0 and 2x - 5 = 0 after solving for x, those would be your answer(s)

carmelle:

ah yes, it would be -2 and 2.5

Angle:

those should be your answers :) what we did was: Step 1: we factored the expression by looking for the two numbers that multiply to the first and third coefficients and added to the middle coefficient Step 2: we split the middle value into the two numbers we found in step 1 Step 3: we factored the two chunks Step 4: we factored the things they had in common Step 5: we set each chunk equal to zero and solved

Angle:

these steps will always work for finding the x-intercept

Angle:

*of a parabola

carmelle:

Thank you so much :)